OpenStudy (anonymous):

Solve the quadratic equation by completing the square. x^2-16x+51=0

OpenStudy (anonymous):

Can someone explain it step by step?

OpenStudy (luffingsails):

Hi... So, the first step is to isolate the terms which have the x variable. So, subtract 51 from each side.

OpenStudy (luffingsails):

Can you try that and write what you get?

OpenStudy (anonymous):

x^2-16x=-51

OpenStudy (luffingsails):

Great... now if there were a coefficient (or number) in front of the x^2 term, we would divide both side by that number. In this case the number is (1)... so no need. Understand?

OpenStudy (anonymous):

Ok, yes.

OpenStudy (luffingsails):

Next step, we look at the coefficient of the x-term (-16), we need to take half of it, square it, and then add it to both sides. Can you try that step and let me know what you get?

OpenStudy (anonymous):

4^2?

OpenStudy (luffingsails):

Not quite... half of -16 = -16/2 = -8. Then we would need to square that. (-8)^2 = 64. We would add 64 to both sides. Make sense?

OpenStudy (anonymous):

Oh so we're not trying to get 16?

OpenStudy (luffingsails):

No, we are trying to complete the square... which is what it means to take the coefficient, halving it and then squaring the result.

OpenStudy (anonymous):

so it would look like this?: (x-8)^2=-64

OpenStudy (luffingsails):

So, what you will have as a result is: x^2 - 16x + 64 = -51 + 64

OpenStudy (luffingsails):

Almost! Just don't forget the first step where we subtracted that 51.

OpenStudy (luffingsails):

Instead, we'll have: (x-8)^2 = 13

OpenStudy (luffingsails):

And of course, we added 64 to each side. <smile>

OpenStudy (luffingsails):

Can you solve it from there?

OpenStudy (anonymous):

x=77 ?

OpenStudy (luffingsails):

Not quite... So, to solve, we first take the square root of each side. x-8 = +/-sqrt(13) x = 8+sqrt(13) or 8 - sqrt(13).

OpenStudy (anonymous):

11.6 and 4.39

OpenStudy (anonymous):

well 11.61

OpenStudy (anonymous):

@luffingsails Sorry to bother you again but I'm practicing on this problem again and I had a question. How did we get 13 in (x-8)^2=13 ??